THE NCAR CSM FLUX COUPLER

by

Frank O. Bryan, Brian G. Kauffman, William G. Large, and Peter R. Gent

National Center for Atmospheric Research ^{1}

P.O. Box 3000 Boulder, CO 80307 U.S.A.

May 1996

1 The National Center for Atmospheric Research is sponsored by the National Science Foundation.

Table of Contents

Pages

A. Introduction . . . . . . . . . . . . . . . . . . . . . . A1–A2 B. The Control Function . . . . . . . . . . . . . . . . . . . B1–B2 C. List of the Information Exchanged . . . . . . . . . . . . C1–C6 D. Temporal Coordination of the Fluxes . . . . . . . . . . . D1–D2 E. Spatial Coordination of the Fluxes . . . . . . . . . . . . . E1–E4 F. Flux Formulae . . . . . . . . . . . . . . . . . . . . . . F1–F6 G. Exceptions and Special Cases . . . . . . . . . . . . . . G1–G2 H. Pseudo-Code for the Flux Coupler . . . . . . . . . . . . H1–H3 I. User’s Guide to the Flux Coupler . . . . . . . . . . . . . I1–I11 J. Input Parameters for the Flux Coupler . . . . . . . . . . . J1–J8 K. Values of Physical Constants . . . . . . . . . . . . . . . K1–K2